find "y'" for each:

1. "y= x\u00b3 sin x"

"y= x\u00b2 + 2x cos x"2. "y= x\u00b2 + 2x" "cos x"

3."y = x\/( sec x + 1)"

"y= x\u00b2 + 2x cos x"

1
Expert's answer
2021-02-02T01:32:05-0500

1) "y'=3x^2\sin x+x^3\cos x"

2) "y'=2x+2\cos x-2x\sin x"

3) "y'=\frac{\sec x +1-x\tan x\sec x}{(\sec x+1)^2}"